OpenStudy (kkutie7):

Calculate the pressure exerted by 1.00 mol of ethene behaving as (100 cm3 = 0.1 L) a) an ideal gas #1: at 273.15 K in 22.414 L #2: at 1000 K in 100 cm3

OpenStudy (kkutie7):

So this is kind of a stupid question because i know i should use Pv=nRT but the "behaving as" part has me confused because none of my other problems state this...

OpenStudy (kkutie7):

also: b) a van der Waals gas #3: at 273.15 K in 22.414 L #4: at 1000 K in 100 cm3 The van der Waals parameters of ethene are: a = 4.471 L2 atm/mol2, b = 0.05714 L/mol.

OpenStudy (photon336):

the first part involves re-arranging the formula for what you need. in this case, the variable you need is pressure. \[IDEAL~GAS |~~~ pV = nRT \] \[\frac{ nRT }{ V } = p\]

OpenStudy (photon336):

yeah, "behaving as" is actually very important

OpenStudy (photon336):

because that will determine what equation you use

OpenStudy (photon336):

if it's an ideal gas you use pV = nRT if it is a vanderwalls gas you will need to use the modified form of the ideal gas equation.

OpenStudy (kkutie7):

so i just used pv=nrt for a and then van der waals for the b (i don't know this one)

OpenStudy (photon336):

yeah so you know that the ideal gas law doesn't really apply to gases in the real world ?

OpenStudy (kkutie7):

yeah i know that i just don't know the van der waals equation. it something with p=rt/vm...

OpenStudy (photon336):

let me explain it to you the first assumption is that gases have no attractions or repulsions. that's false. the reason is that gases are always bumping into the walls of the container and sometimes colliding with each-other. now let's try to define pressure that's force per area. \[p = \frac{ F }{ A }\] so we need to account for this. another ideal assumption is that gases occupy zero volume. this is not really the case because gases take up space so in actuality the volume that the gases have to move around is less. the volume available will be the volume of the container minus the volume that the gas takes up i.e the number of moles of gas times some constant. that's how we get (v-nB) now we put both of these together and we get the formula below. \[(p+\frac{ n^{2}a^{2} }{ v })(V-n~\beta) = nRT \] solving this for pressure is easy, but i believe that alpha and beta would depend on the gas you have. \[p+\frac{ n^{2}\alpha }{ V } = \frac{ nRT }{ (V-n \beta) }\]

OpenStudy (kkutie7):

The van der Waals parameters of ethene are: a = 4.471 L2 atm/mol2, b = 0.05714 L/mol.

OpenStudy (photon336):

\[P = \frac{ nRT }{ V-n \beta }-\frac{ n^{2} }{ V^{2} }*\alpha\]

OpenStudy (photon336):

the idea now is just to put everything together.

OpenStudy (kkutie7):

ok Im pretty sure i can do that thank you

OpenStudy (photon336):

the attractions and repulsions or alpha and beta will be different for each gas. those are properties that depend on the gas you're talking abot.

OpenStudy (photon336):

the last thing for you to focus on would be to make sure that your units are consistent, then you just need to plug the values in and see if you get the right units for pressure.